One number is 12 more than another. the sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by checkley71(8403)

X=Y+12
X+2(X+12)=39
X+2X+24=39
3X=39-24
3X=15
X=15/3
X=5 THE SMALLER NUMBER.
5+12=17 FOR THE LARGER NUMBER.
PROOF
5+2*17=39
5+34=39
39=39